1. Andrea, a self-employed individual, wishes to accumulate a retirement fund of $250,000. How much should she deposit each month into her retirement account, which pays interest at a rate of 2.5%/year compounded monthly, to reach her goal upon retirement 35 years from now? (Round your answer to the nearest cent.)

2. Joe secured a loan of $13,000 five years ago from a bank for use toward his college expenses. The bank charges interest at the rate of 3%/year compounded monthly on his loan. Now that he has graduated from college, Joe wishes to repay the loan by amortizing it through monthly payments over 11 years at the same interest rate. Find the size of the monthly payments he will be required to make. (Round your answer to the nearest cent.)
$
Answer #1

Future value = pmtx ((1+r)^n-1))/ PMT = FV = ((1+r)^n-1))/r Save each month = PMT= $372.90 250,000/(((1.0020831420)-1)/0.0020